OpenStudy (anonymous):

If T+10 indicates 10 second after takeoff, what interger indicates 10 seond before takeoff? What interger represents takeoff?

jimthompson5910 (jim_thompson5910):

What do you think it is?

OpenStudy (anonymous):

I'm really not sure?

OpenStudy (anonymous):

Oh could it be T-10

jimthompson5910 (jim_thompson5910):

T+10 = 10 seconds after take off T+9 = 9 seconds after take off ...

jimthompson5910 (jim_thompson5910):

yes it would be, T - 10 is 10 seconds before takeoff

jimthompson5910 (jim_thompson5910):

What interger represents takeoff?

jimthompson5910 (jim_thompson5910):

integer*

OpenStudy (anonymous):

I'm not sure

jimthompson5910 (jim_thompson5910):

T+10 = 10 seconds after take off T+9 = 9 seconds after take off T+8 = 8 seconds after take off T+7 = 7 seconds after take off T+6 = 6 seconds after take off T+5 = 5 seconds after take off ... etc etc Keep going until you reach takeoff

OpenStudy (anonymous):

Oh is it T

jimthompson5910 (jim_thompson5910):

yes it is T

OpenStudy (anonymous):

Thank you so much

jimthompson5910 (jim_thompson5910):

yw